FLAO Task List Jan. 26, 2017

  • List of near term AO tasks
  • V0.2
  • 26 Jan. 2017
  • Do not distribute

Email the above distribution additional tasks or modifications. This list will be sent to a larger distribution after few iterations and (possibly) discussion on Thursday

These tasks are in random order (not priority order)

Completion Date: March 6th, 2017, ARGOS run begins

led-yellow WFS pupil shift during telescope offset

  • IT #6333
  • Affects AO nighttime operation: Limits the size of offset to <30"
  • Data Collection
    • We have collected several sets of long AO closed loop telemetry (see IT #6333) and are waiting for analysis from Arcetri
  • Resources
    • Arcetri: Alfio
    • Nighttime: 0.25 nights for testing
  • Complete date: By ARGOS run, March 2, 2017
  • Status: Waiting on Arcetri analysis

led-yellow AO stage transformations on SX and DX

  • Measure and confirm
  • LBTO: Doug/Greg
  • On-sky (0.25 nights)
  • Affects AO nighttime operation: Offsets may not be properly aligned with detector
  • Status: Doug's needs to re-process old transform data for SX and DX

led-yellow Implement Tip/Tilt mirror Lookup Table to correct for rerotator (K-Mirror) misalignment

  • Both SX and DX
  • Data Collection
    • We have already measured a "rough" lookup table using the ARGOS CalUnit. Thus, the next step is to develop and build a process that will automatically apply this LUT offset voltage to the Tip/Tilt mirror
    • We will probably have to collect more measurement to refine the LUT. This can be done with the ARGOS CalUnit.
    • If the IRTC is mounted on the DX side when LUCI2 is in the lab for repair, we can use the IRTC for daytime measurements. In fact, it will be much easier to remeasure the LUT with the IRTC because we and rotate the instrument rotator and not have to worry or correct for instrument flexure. IRTC has been measure to have no flexure due to rotation.
  • Resources
    • LBTO: Doug, data collection, testing
    • LBTO: Xianyu, algorithm construction
    • Arcetri: Alfio, advise
    • Daytime: ~1 day of Doug and Xianyu's time for algorithm building and testing
    • Daytime: 1 day with ARGOS CalUnit for data collection
    • Daytime: 1 day with ARGOS CalUnit for testing
    • Nighttime: 0.25 nights for testing
  • Status: Waiting for advise from Arcetri
    • Doug pinged Alfio yesterday. They now have time to look into it

led-red Purchase four stand alone IDL licenses for the mountain AdSec and WFS machines

  • Doug and Xianyu will purchase and install

led-green CCD39 noise problem Repair

  • IT #6319
  • LBTO: Gustavo/Alessandro
  • 1 Week in Mountain lab
  • Affects AO nighttime operation: decreases AO performance for all magnitudes

led-green CCD39 fix Test

  • LBTO: Greg/Doug
  • ARGOS CalUnit (0.5 days)
  • On-sky (0.25 nights)

Completion Date: March 14th, 2017, LUCI/AO commissioning begins

led-red Develop and include automatic starting of Gopt/NCPA

  • LBTO: Doug/Greg/Xianyu + Alfio support
  • Weeks of development
  • On-sky testing (0.25 nights)

led-green Insert new NCPA amplitudes into SX and DX configurations files

  • LBTO Greg
  • ~ hour+

led-red Polynomial fit new NCPA amplitude values

  • Repeat fits performed by Doug on original data
  • Create and install new NCPA lookup table from fit values
  • LBTO: Doug
  • 0.5 days

Completion Date: TBD

led-red \x93unknown exception from arbitrator\x94 error with UAO

  • IT #6353
  • Arcetri: Alfio
  • Days of analysis/repair
  • Affects AO nighttime operation: Lost time if the error occurs ~ 20 minutes

led-green AO Supervisor configuration in new LBTO svn

  • Migrate modified configuration files from ROOT to SOURCE directory
  • LBTO: Doug
  • 0.5 days

led-green AO Supervisor from LBTO svn test

  • LBTO: Doug/Greg/Xianyu
  • ARGOS CalUnit (1.0 day)
  • On-sky (0.25 nights)

led-yellow Continued review of \x93AO Operations\x94 web pages

  • LBTO: AO Group
  • Status: Continuing review

led-red AO Operations training for TO\x92s

  • LBTO: Doug/Greg
  • On-sky (parallel to AO observations)

led-red ELAB improvements

  • Include Gopt scale factor in calculation of modalplot, etc
  • Understand/fix telemetry data files
  • Add automatic telemetry data collection/analysis (?)
    • LBTO: Greg/Doug/Xianyu/Pedro + Alfio support
    • Weeks of development

led-yellow Log Crawler development

  • LBTO: Doug
  • Weeks/months of development
  • Status: Doug works on this as time allows

led-red Mirror of AO data in Tucson

  • AO logs mirror
  • AO Telemetry mirror
    • LBTO: Xianyu/IT/Doug/Greg

led-red Develop semi-automatic windowing of LUCI sub window

  • LBTO: Doug/Greg/Xianyu/Tom + LUCI support
  • Days/weeks of development
  • ARGOS CalUnit (1.0 days)

led-yellow Develop AO test scripts to check AO system closed dome at handover

  • LBTO: Greg + Alfio support
  • Days
  • Status: Greg has some scripts ready for testing use. More to come

led-red Develop AO plan and scripts for start of night performance check

  • LBTO: Greg/Doug
  • Days
  • on-sky beginning of the 5 nights

led-red Implement the monitoring script functionality to UAO

  • LBTO: Xianyu

led-red Additional safety ASM action from software

  • What is this?

led-red Transfer the AdSec safety information to TCS/Alarm handler through AOS

  • Pass information via AOS to DD variable instead of current "back door"
  • LBTO: Xianyu

led-red ASM dump file analysis tool

  • LBTO

led-red Add monitor routines to WUnit (similar to AdSec monitor routines)

  • LBTO: Xianyu

led-red Zabbix on WFS computers

  • Zabbix move to CentOS 6 machines
  • mysql moved to CentOS 6 machines
  • LBTO: Xianyu + support from Arcetri

led-red New AO State machine GUI

* Intervention mode with options * Need ICE interface commands to AO Supervisor * LBTO: Doug/Xianyu

led-yellow Create GUI to track AO daytime testing

  • Either web based or python based)
  • Filled out by whomever is going to do the testing
  • Automatically written to a TWiki page (like Nightly Reports)
  • Send email to telescope work and others
  • Greg receive information from Steve A.
  • Status: Greg will look at Steve routines as possible starting framework

led-yellow Create GUI to record AO on-call activities

  • Filled out by AO Support person that performs the work
  • Automatically written to a TWiki page (like Nightly Reports)
  • Send email to AOgroup and others
  • Greg receive information from Steve A.
  • Status: Greg will look at Steve routines as possible starting framework

-- %USERSIG{DougMiller - 2017-01-26}%
Topic revision: r7 - 17 Feb 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback